# Squish Jamming

### v1

@@ III Large geometrical deformations absorb shear and hinder rearrangements
We now focus on the deformation of rings during shear cycles to understand how yielded and jammed regions respond to shear. The geometry, of the rings is quantified by the shape index $q=(\mathrm{ring~{}perimeter})/\sqrt{\mathrm{ring~{}area}}$, a classic quantity used in highly deformable assemblies like epithelial tissues [16, 17] to assess how far a convex shape is from a circle. The shape index takes the value $q=2\sqrt{\pi}\approx 3.54$ for a perfect circle and increases when the ring gets deformed. As the rings are slender objects, they deform mostly by flexion rather than stretching. Thus, $q$ is mainly increased by reducing the area while the perimeter is almost unchanged. This contrasts with emulsions or tissues where most deformations originate from a change in perimeter and not the area [7, 27]. The perimeter and area of the rings are obtained via image analysis, and a Fourier decomposition is used to smooth the pixelated shape and reduce the shape to a few descriptors [27]. The resulting measure of $q$ during a shear cycle is shown in Fig. 3a for one ring within an assembly with parameters $(\gamma_{0},\Phi)=(0.4,1.00)$. For each cycle $c$ and tracked ring, we define two quantities from the shape trajectory: the average deformation $\tilde{q}=\left\langle{q}\right\rangle_{\mathrm{cycle}}$ and the geometrical fluctuation $\Delta q=\mathrm{max}(q)_{\mathrm{cycle}}-\mathrm{min}(q)_{\mathrm{cycle}}$.
@@
In the steady state, the mean shape index and fluctuation are then averaged over all tracked rings, and their dependence on shear is compared for each density Fig. 3b,c. As expected, the average geometrical deformation $\tilde{q}$ increases steadily with the density Fig. 3b. For all densities, increasing shear amplitude drives significant shape fluctuations during shear cycles, and the higher the density, the more pronounced the geometrical changes (Fig. 3c). Thus, a higher density contributes, in addition to naturally increasing the average geometrical deformation, to promoting shape changes for a given imposed shear. Ring assemblies, therefore, use their compressibility and shape-changing abilities to absorb imposed external shear, and they do so more efficiently as the density increases.
@@
To grasp the combined role of geometrical changes and structural rearrangements leading to jammed or yielded assemblies, we quantify these two aspects by dedicated strains derived for epithelial tissue deformations [15]. For each ring and between each time frame, we track the bond vectors between touching neighbours (Fig. 1e and Supplementary Video 6). The geometrical strain $S^{*}$ is derived from the change in bond length and orientation and quantifies changes in size and shape within the assembly. In contrast, the topological strain $T^{*}$, based on the appearance or creation of bonds due to rearrangements, captures the change in the topology of the contact network and then quantifies the structural deformations of the assembly. More details on these quantities are given in the Methods, and their step-by-step derivation is in the Appendix of Ref.[15]. The strains $S^{*}$ and $T^{*}$ are computed between each time frame and for the whole assembly. Their norm is then summed over all the cycles in the steady states to obtain the quantities $S_{c}=\int_{t}\norm{S^{*}}\mathrm{d}t$ and $T_{c}=\int_{t}\norm{T^{*}}\mathrm{d}t$ shown in Fig. 3d,e. From the map of $\log S_{c}$ over $(\gamma_{0},\Phi)$ in Fig. 3d, we observe that geometrical changes increase significantly with both shear amplitude and density as observed in Fig. 3c. By contrast, the map of $\log T_{c}$ in Fig. 3e shows that structural deformations, while increasing with shear amplitude, are diminished by higher densities. As a result, the region where changes in the topology of the contact network are significant coincides with the fluid region identified in Fig. 1g,h. Yet, the geometrical deformations vary regardless of the fluid or solid regions.
In contrast to confluent bi-dimensional active cell models where more significant deformations lead to yielding [18], here, deforming the assembly by increasing the density fosters jamming by absorbing shear and hindering structural rearrangements. We then investigate the ring-ring interactions to understand the mechanisms under which large shape deformations impede structural ones.
@@
IV Geometry and friction concur to generate effective shear adhesion between rings.
@@
Figure 4: Friction and geometrical deformations produce an effective shear adhesion between the rings. a, Ring-ring interaction experiments and FEM simulations. Two facing half-rings are compressed over a distance $\delta_{y}$ and then sheared laterally until they lose contact. Displayed are two FEM snapshots showing the contact pressure between the rings after compression, before (top) and after (bottom) a lateral shear displacement of $x=r$. See also Supplementary Video 7. b,c, The resulting shear interaction force $f_{x}$ is measured as a function of the lateral displacement $x$. Negative (positive) values indicate an attractive (repulsive) force between the rings. Experimental responses are averaged over $3$ distinct pairs, and shaded areas represent max and min values obtained from the different pairs. Panel b shows the shear response $f_{x}(x)$ for $\delta_{y}=950\,\mathrm{\mu m}$ and the friction coefficient $\mu=0.1,0.3,0.6$ varied in the FEM. The experimental response matches that of FEM for $\mu=0.6$. Panel c shows the shear response $f_{x}(x)$ for $\mu=0.6$, for increasing confinement $\delta_{y}=200,600,950\,\mathrm{\mu m}$. In this high friction regime, $f_{x}$ remains negative until the separation of the rings. We measure $x_{r}(\delta_{x},\mu)$, the displacement needed to switch from attractive ($f_{x}<0$) to repulsive ($f_{x}>0$) shear interactions. d, Experimental measure of the attractive-repulsive threshold shear $\gamma_{r}(\Phi)$, obtained by translating local quantities $(\delta_{y},x_{r})$ to global ones $(\Phi,\gamma_{r})$, together with results from FEM (dashed lines) and the geometrical predictions (dotted lines) for $\mu=0.1,0.3,0.6$. This curve $\gamma_{r}(\Phi)$ for $\mu=0.6$ is the one drawn in the phase diagrams shown so far (e.g., Fig. 1g,h), marking the fluid-solid transition of the assembly (fluid-like for $\gamma_{0}>\gamma_{r}$ and solid-like otherwise). e, $(\Phi,\gamma_{0})$ map of the difference between the imposed shear displacement $x_{0}$ and the measured maximal contact size $\langle{\ell_{c}^{\mathrm{max}}}\rangle$, averaged over all tracked rings in the steady state. The geometrical prediction and local measurements of the attractive-repulsive threshold $\gamma_{r}$ correspond to the region where $y_{0}\approx\langle{\ell_{c}^{\mathrm{max}}}\rangle$. In this high friction and large geometrical deformations limit, the extent of the contacts $\ell_{c}$ mainly controls the jamming transition.
@@
We introduce a microscopic ersatz comprising two half-rings sequentially compressed and sheared to measure the interaction between rings. Then, we will link the sign of the shear force, which would promote or prevent structural rearrangements, to the assembly’s fluid-solid transition.
@@
We measure, via FEM and experiments, the shear response of two clamped rings compressed by an amount $\delta_{y}$ (Fig. 4a and Supplementary Video 7). With FEM, we systematically vary the frictional interaction via the friction coefficient $\mu$ of the sheared interfaces (Fig. 4b) and the compression $\delta_{y}$ between the rings (Fig. 4c). Experimentally, we test rings taken from the assembly by applying the same compression values as in FEM. We find the best agreement between the experiments and the FEM for $\mu=0.6$, constituting a measure of the experimental friction coefficient. In the absence of friction, the shear force $f_{x}$ is instantly positive (repulsive) as elastic deformations push the rings away from each other. With friction, however, $f_{x}$ starts negative (attractive) and a finite lateral displacement $x_{r}(\delta_{y},\mu)$ is required to make it repulsive ($f_{x}>0$). The synergy of geometrical deformation and friction thus produces an attractive interaction, or effective adhesion, between the rings until sufficient shear is applied. The role of shape change to generate an effective adhesion under shear is reminiscent of the rheology of cohesive granular media [49], where the low stiffness of the grains also participates in producing a higher effective adhesion. We argue next that the presence or absence of this effective adhesion under shear is at the root of the assembly’s fluid or solid macroscopic response.
@@
We now give a geometrical prediction of the finite lateral displacement $x_{r}$ necessary to switch the shear interactions force from attractive to repulsive, i.e., $f_{x}(x_{r})=0$. This distance is a function of the initial compression $\delta_{y}$ and the friction between the rings, $\mu$. Two cases emerge in the way $x_{r}(\delta_{y},\mu)$ is determined. In the first case, the rings remain in contact when $f_{x}$ switches from negative to positive (e.g., Fig. 4b for $\mu=0.1$ and $0.3$). In this case, assuming Amontons-Coulomb frictional properties and reducing the contacting interfaces to a point, one can directly link the sign of the shear force $f_{x}$ to the lateral displacement and obtain $x_{r}(\delta_{y},\mu)=\mu(2r_{h}-\delta_{y})$. By contrast, in the second case where $f_{x}$ remains negative until the rings are fully separated (e.g., Fig. 4c), $x_{r}$ is given by the extent of the initial overlap between the rings: $x_{r}(\delta_{y},\mu)=2r_{h}\sqrt{1-(1-\delta_{y}/(2r_{h}))^{2}}$. Now that we have a prediction for the microscopic ring-ring interactions, we must translate the compression $\delta_{y}$ and local lateral displacement $x$ to the global variable density $\Phi$ and shear $\gamma$. Assuming an hexagonal structure of density $\Phi_{h}$, an isotropic local overlap $\delta_{y}$ directly translates into a density by $\Phi(\delta_{y})=\Phi_{h}/(1-\delta_{y}/(2r_{h}))^{2}$. For the shear $\gamma$, too, by using its local counterpart, i.e., the local shear imposed between the two half rings, we obtain $\gamma=x/(2r_{h}-\delta_{y})$. Schematics of the underlying geometrical predictions are shown in Extended Data Fig. 2. By combining the expressions of $x_{r}(\delta_{y},\mu)$ with the expressions of $\Phi(\delta_{y})$ and $\gamma(y,\delta_{y})$, we can finally express the repulsive shear $\gamma_{r}$ at which the interactions become repulsive as a function of global variables $(\gamma,\Phi)$, as follows:
@@
In Fig. 4d, we compare $\gamma_{r}$ from the prediction against FEM and experimental measurements and find a perfect match. In the high friction case ($\gamma_{r}<\mu$), the threshold $\gamma_{r}(\Phi)$ is simply set by the extent of contacts, yielding $\gamma_{r}(\Phi)=\sqrt{\Phi/\Phi_{h}-1}$. Experimentally, the explored density range and the estimated friction coefficient $\mu\approx 0.6$ are such that we remain in this high friction limit. In soft spheres without shape changes, the jamming transition also features a scaling between yielding and density [50] but occurs at a much lower density and with a different scaling.
@@
With repulsive interactions between rings, we expect structural rearrangements to be promoted and, inversely, hindered if interactions are attractive and act as an effective adhesive force. The shear $\gamma_{r}$ can then be viewed as a yield shear, where rearrangements leading to fluid-like response are expected for $\gamma_{0}>\gamma_{r}$, and a solid-like jammed assembly for $\gamma_{0}<\gamma_{r}$. Indeed, the line drawn by the geometric prediction of Eq. (1) qualitatively separates the fluid and solid-like regions in the $(\Phi,\gamma_{0})$ parameter space (Fig. 1g,h), which also accompany significant changes of the contact network topology (Fig. 3e). This geometric boundary $\gamma_{r}(\Phi)=\sqrt{\Phi/\Phi_{h}-1}$ is set by the extent of the contact between rings, which increases with density. To verify that the contact extent is a major factor controlling the fluid-solid transition, we measure ${\ell_{c}^{\mathrm{max}}}$, the largest contact length among all touching neighbours, for each ring and at each time step. We then compute the difference between the equivalent local lateral displacement amplitude $x_{0}=\gamma_{0}(2r_{h}-\delta_{y})=\gamma_{0}2r_{h}\sqrt{\Phi_{h}/\Phi}$ and $\langle{\ell_{c}^{\mathrm{max}}}\rangle$, averaged in the steady state (Fig. 4e). The geometrical prediction of $\gamma_{r}$ matches with the region where $x_{0}\approx\langle{\ell_{c}}^{\mathrm{max}}\rangle$. The fluid-solid regions are then defined by whether the shear is large enough to overcome the effective adhesion and move the rings by a distance greater than the maximal contact length. Therefore, in contrast with rigid granular media where the number of contacts controls the jamming transition [4], the transition with squishy grains in a high friction limit seems to be strongly controlled by the extent of the contacts (Extended Data Fig. 3).
V Discussion

### v2

@@ Abstract
We now focus on the deformation of rings during shear cycles to understand how the yielded and rigid regions respond to shear. The geometry of the rings is quantified by the shape index $q=(\mathrm{ring\leavevmode\nobreak\ perimeter})/\sqrt{\mathrm{ring\leavevmode\nobreak\ area}}$, a classic quantity used in highly deformable assemblies like epithelial tissues (17, 18) to assess how far a convex shape is from a circle. The shape index takes the value $q=2\sqrt{\pi}\approx 3.54$ for a perfect circle and increases when the ring gets deformed. Since the rings are slender objects, they deform mostly by flexion rather than stretching. Thus, $q$ is mainly increased by reducing the area while the perimeter is almost unchanged. This contrasts with emulsions or tissues where most deformations originate from a change in perimeter and not the area (7, 29). The perimeter and area of the rings are obtained via image analysis, and a Fourier decomposition is used to smooth the pixelated shape and reduce the shape to a few descriptors (29). The resulting measure of $q$ during a shear cycle is illustrated in Fig. 4A for one arbitrary ring within an assembly with parameters $(\gamma_{0},\Phi)=(0.4,1.00)$. The diversity of shape trajectories for all the rings for a given cycle (Fig. 4B) results from the heterogeneity of the assembly, but nonetheless the averaged trajectory is symmetric within a shear cycle. For each cycle $c$ and tracked ring, we define two quantities from the shape trajectory: the average shape index $\tilde{q}=\left\langle{q}\right\rangle_{\mathrm{cycle}}$ and the variation amplitude $\Delta q=\mathrm{max}(q)_{\mathrm{cycle}}-\mathrm{min}(q)_{\mathrm{cycle}}$ (Fig. 4A), quantifying the mean deformation and the geometrical fluctuation, respectively.
@@
In the steady state, the mean shape index and fluctuation are then averaged over all tracked rings, and their dependence on shear is compared for each density (Fig. 4C,D). As expected when squeezing more rings in the same volume, the average geometrical deformation $\langle\tilde{q}\rangle$ increases steadily with the density (Fig. 4C). While increasing shear amplitude has limited influence on $\langle\tilde{q}\rangle$, it drives significant shape fluctuations during shear cycles for all densities (Fig. 4D). Moreover, the higher the density, the more pronounced the geometrical changes. Thus, a higher density contributes, in addition to naturally increasing the average geometrical deformation, to promoting shape changes for a given imposed shear. Ring assemblies, therefore, use their compressibility and shape-changing abilities to absorb imposed external shear, and they do so more efficiently as the density increases.
@@
To grasp the combined role of geometrical changes and structural rearrangements leading to solid or yielded assemblies, we quantify these two aspects by dedicated strains used previously for epithelial tissue deformations (16), namely the geometrical strain $S^{*}$ and the topological strain $T^{*}$. These can be evaluated from the dynamics of the bond vectors connecting touching neighbors (Fig. 1D and Movie S6). While the precise definition is given in Methods (see also Ref.(16)), essentially the geometrical strain $S^{*}$ is a contribution from the change in bond length and orientation and quantifies solely the continuous change in the size and shape of the bond network. In contrast, the topological strain $T^{*}$ is based on the creation and annihilation of bonds due to rearrangements and captures the discontinuous, topological change in the contact network. The strains $S^{*}$ and $T^{*}$ are computed between each time frame and for the whole assembly. Their norm is then summed over all the cycles in the steady states to obtain the quantities $S_{c}=\int_{t}\norm{S^{*}}\mathrm{d}t$ and $T_{c}=\int_{t}\norm{T^{*}}\mathrm{d}t$ shown in Fig. 4E,F, which quantify the degree of the geometrical deformation and the topological rearrangements, respectively, of the ring assembly over the shear cycles. The map of $\log S_{c}$ in the parameter space $(\gamma_{0},\Phi)$ (Fig. 4E) reveals that geometrical changes increase significantly with both shear amplitude and density, regardless of the yielded or solid regions. By sharp contrast, the map of $\log T_{c}$ in Fig. 4F shows that topological deformations, while increasing with shear amplitude, are diminished for higher densities. As a result, the region with significant changes in the topology of the contact network coincides with the yielded region identified in Fig. 2C and $D$. Therefore, while the yielded region features both topological rearrangements and geometrical deformations, the solid region is characterized by the quasi-absence of topological rearrangements, and consequently, the imposed shear is entirely absorbed by the geometrical deformations of the rings. It then appears that, for a fixed shear amplitude, an increase in density (and hence more deformed assemblies) hinders the structural rearrangements, leading to the solidification of the assembly. This result contrasts with confluent bi-dimensional active cell models, where more significant deformations lead to yielding (19). Now, to understand how large shape deformations impede structural ones in our ring assembly, we investigate the ring-ring interactions.
Figure 4: Competition between geometrical and structural deformations. (A) The geometrical deformation of the rings is quantified by the shape index $q$. The evolution of $q$ within a cycle is shown for one ring in a $(\gamma_{0},\Phi)=(0.4,1.00)$ assembly. All along the shape trajectory, the shape of the ring at each diamond marker is shown in the snapshot above it. For each ring and each cycle, we extract the time-averaged shape index $\tilde{q}$ and the variation amplitude $\Delta q$ of its shape trajectory. (B) For the same assembly $(\gamma_{0},\Phi)=(0.4,1.00)$, the shape trajectories of all the rings are represented by light red lines, while the average trajectory is marked by a black line. (C,D) The averages $\langle\tilde{q}\rangle$ (C) and $\langle\Delta q\rangle$ (D) evaluated over all tracked rings and in the steady state. (E,F) Map of the geometrical $S_{c}$ and topological $T_{c}$ strains of the assembly. Strains are estimated using the method developed in (16) (see also Methods) based on the frame-to-frame tracking of bond vectors between touching neighboring rings. The total strains here are the cumulative sums of the norm of the frame-to-frame strains estimated over the whole assembly and in the steady state.
@@
Geometry and friction concur to generate effective shear adhesion between rings
@@
Figure 5: Friction and geometrical deformations produce an effective shear adhesion between the rings. (A) FEM simulations and experiments of ring-ring interactions. Two facing half-rings are compressed over a distance $\delta_{y}$ and then sheared laterally until they lose contact. The FEM snapshot shows the contact pressure between the rings after compression and a lateral shear displacement of $x$. See also Movie S7. (B,C) The resulting shear interaction force $f_{x}$ is measured as a function of the lateral displacement $x$. Negative (positive) values indicate an attractive (repulsive) force between the rings. Experimental responses are averaged over $3$ distinct pairs, and shaded areas represent max and min values obtained from the different pairs. Panel (B) shows the shear response $f_{x}(x)$ for $\delta_{y}=950\,\mathrm{\mu m}$ and the friction coefficient $\mu=0.1,0.3,0.6$ varied in the FEM. The experimental response matches that of FEM for $\mu=0.6$. Panel (C) shows the shear response $f_{x}(x)$ for $\mu=0.6$, and increasing confinement $\delta_{y}=200,600,950\,\mathrm{\mu m}$. In this high friction regime, $f_{x}$ remains negative until the separation of the rings. We measure $x_{r}(\delta_{x},\mu)$, the displacement needed to switch from attractive ($f_{x}<0$) to repulsive ($f_{x}>0$) shear interactions. (D) Schematic of the ring-ring interaction on which the geometrical predictions are based. At a given shear distance $x$, the resulting shear force $f_{x}$ can either take positive or negative values depending on the friction coefficient (yellow and red arrows, respectively). (E) Experimental measure of the attractive-repulsive threshold shear $\gamma_{r}(\Phi)$, obtained by translating local quantities $(\delta_{y},x_{r})$ to global ones $(\Phi,\gamma_{r})$, together with results from FEM (dashed lines) and the geometrical predictions (dotted lines) for $\mu=0.1,0.3,0.6$. This curve $\gamma_{r}(\Phi)$ for $\mu=0.6$ is the one drawn in the phase diagrams shown so far (e.g., Fig. 2C and $D$), marking the rigidity transition of the assembly (yielded for $\gamma_{0}>\gamma_{r}$ and solid-like otherwise). (F) $(\Phi,\gamma_{0})$ map of the difference between the imposed shear displacement $x_{0}$ and the measured maximal contact size $\langle{\ell_{c}^{\mathrm{max}}}\rangle$, averaged over all tracked rings and in the steady state. The geometrical prediction and local measurements of the attractive-repulsive threshold $\gamma_{r}$ correspond to the region where $x_{0}\approx\langle{\ell_{c}^{\mathrm{max}}}\rangle$. In this high friction and large geometrical deformations limit, the extent of the contacts $\ell_{c}$ mainly controls the rigidity transition.
@@
We introduce a microscopic ersatz comprising two half-rings sequentially compressed and sheared to measure the interaction between rings. Then, we will link the sign of the shear force, which would promote or prevent structural rearrangements, to the assembly’s rigidity transition.
@@
We measure, via FEM and experiments, the shear response of two clamped rings compressed by an amount $\delta_{y}$ (Fig. 5A and Movie S7). With FEM, we systematically vary the frictional interaction via the friction coefficient $\mu$ of the sheared interfaces (Fig. 5B) and the compression $\delta_{y}$ between the rings (Fig. 5C). Experimentally, we test rings taken from the assembly by applying the same compression values as in FEM. We find the best agreement between the experiments and the FEM for $\mu=0.6$, constituting a measure of the experimental friction coefficient. In the absence of friction, the shear force $f_{x}$ would be instantly positive (repulsive) as elastic deformations push the rings away from each other. With friction, however, $f_{x}$ starts negative (attractive) and a finite lateral displacement $x_{r}(\delta_{y},\mu)$ is required to make it repulsive ($f_{x}>0$) (see Fig. 5B). The synergy of geometrical deformation and friction thus produces an attractive interaction, or effective adhesion, between the rings until sufficient shear is applied. The role of shape change to generate an effective adhesion under shear is reminiscent of the rheology of cohesive granular media (55), where the low stiffness of the grains also participates in producing a higher effective adhesion. We argue next that the presence or absence of this effective adhesion under shear is at the root of the assembly’s yielded or solid macroscopic response.
@@
We now give a geometrical prediction of the finite lateral displacement $x_{r}$ necessary to switch the shear interaction force from attractive to repulsive, i.e., $f_{x}(x_{r})=0$. This distance is a function of the initial compression $\delta_{y}$ and the friction between the rings, $\mu$. Two cases emerge in the way $x_{r}(\delta_{y},\mu)$ is determined (Fig. 5D). In the first case, the rings remain in contact when $f_{x}$ switches from negative to positive (e.g., Fig. 5B for $\mu=0.1$ and $0.3$). In this case, assuming Amontons-Coulomb frictional properties and reducing the contacting interfaces to a point, one can directly link the sign of the shear force $f_{x}$ to the lateral displacement and obtain $x_{r}(\delta_{y},\mu)=\mu(2r_{h}-\delta_{y})$. By contrast, in the second case where $f_{x}$ remains negative until the rings are fully separated (e.g., Fig. 5C), $x_{r}$ is given by the extent of the initial overlap between the rings: $x_{r}(\delta_{y},\mu)=2r_{h}\sqrt{1-(1-\delta_{y}/(2r_{h}))^{2}}$. Now that we have a prediction for the microscopic ring-ring interaction, we must translate the compression $\delta_{y}$ and local lateral displacement $x$ to the global variable density $\Phi$ and shear $\gamma$. Assuming an hexagonal structure of density $\Phi_{h}$, an isotropic local overlap $\delta_{y}$ directly translates into a density by $\Phi(\delta_{y})=\Phi_{h}/(1-\delta_{y}/(2r_{h}))^{2}$. For the shear $\gamma$, too, by using its local counterpart, i.e., the local shear imposed between the two half rings, we obtain $\gamma=x/(2r_{h}-\delta_{y})$. Schematics of the underlying geometrical predictions are shown in SI Appendix, Fig. LABEL:S-fig_sup_schem. By combining the expressions of $x_{r}(\delta_{y},\mu)$ with the expressions of $\Phi(\delta_{y})$ and $\gamma(y,\delta_{y})$, we can finally express the repulsive shear $\gamma_{r}$ at which the interactions become repulsive as a function of global variables $(\gamma,\Phi)$, as follows:
@@
In Fig. 5E, we compare $\gamma_{r}$ from the prediction against FEM and experimental measurements and find a perfect match. In the high friction case ($\gamma_{r}<\mu$), the threshold $\gamma_{r}(\Phi)$ is simply set by the extent of contacts, yielding $\gamma_{r}(\Phi)=\sqrt{\Phi/\Phi_{h}-1}$. Experimentally, the explored density range and the estimated friction coefficient $\mu\approx 0.6$ are such that we remain in this high friction limit. In soft spheres without shape changes, the jamming transition also features a scaling between yielding and density (56) but occurs at a much lower density and with a different scaling.
@@
With repulsive interactions between rings, we expect structural rearrangements to be promoted and, inversely, hindered if interactions are attractive and act as an effective adhesive force. The shear $\gamma_{r}$ can then be viewed as a yield shear, where rearrangements leading to a yielded assembly are expected for $\gamma_{0}>\gamma_{r}$, and a solid-like assembly for $\gamma_{0}<\gamma_{r}$. Indeed, the line drawn by the geometric prediction of Eq. [1] qualitatively separates the yielded and solid-like regions in the $(\Phi,\gamma_{0})$ parameter space (Fig. 2C and $D$), which also accompany significant changes of the contact network topology (Fig. 4F). This geometric boundary $\gamma_{r}(\Phi)=\sqrt{\Phi/\Phi_{h}-1}$ is set by the extent of the contact between rings, which increases with density. To verify that the contact extent is a major factor controlling the rigidity transition, we measure ${\ell_{c}^{\mathrm{max}}}$, the largest contact length among all touching neighbors, for each ring and at each time step. We then compute the difference between the equivalent local lateral displacement amplitude $x_{0}=\gamma_{0}(2r_{h}-\delta_{y})=\gamma_{0}2r_{h}\sqrt{\Phi_{h}/\Phi}$ and $\langle{\ell_{c}^{\mathrm{max}}}\rangle$, averaged in the steady state (Fig. 5F). The geometrical prediction of $\gamma_{r}$ matches with the region where $x_{0}\approx\langle{\ell_{c}}^{\mathrm{max}}\rangle$. The yielded and solid regions are then defined by whether the shear is large enough to overcome the effective adhesion and move the rings by a distance greater than the maximal contact length. Therefore, in contrast with rigid granular media where the number of contacts controls the jamming transition (4), the transition with squishy grains in a high friction limit seems to be strongly controlled by the extent of the contacts (SI Appendix, Fig. LABEL:S-fig_sup_coor).
Discussion